# Millinovae: a new class of transient supersoft X-ray sources without a nova eruption

### v1

@@ Abstract
Some accreting binary systems containing a white dwarf (such as classical novae or persistent supersoft sources) are seen to emit low energy X-rays with temperatures of 106 K and luminosities exceeding erg s-1. These X-rays are thought to originate from nuclear burning on the white dwarf surface, either caused by a thermonuclear runaway (classical novae) or a high mass accretion rate that sustains steady nuclear burning (persistent sources). The discovery of transient supersoft X-rays from ASASSN-16oh challenged these ideas, as no signatures of nuclear fusion were detected, and the origin of these X-rays remains controversial. It was unclear whether this star was one of a kind or representative of a larger, as yet undiscovered, group. Here we present the discovery of 29 stars located in the direction of the Magellanic Clouds exhibiting long-duration, symmetrical optical outbursts similar to that seen in ASASSN-16oh. We observed one of these objects during an optical outburst and found it to be emitting transient supersoft X-rays, while no signatures of mass ejection (indicative of a classical nova eruption) were detected. We therefore propose that these objects form a homogeneous group of transient supersoft X-ray sources, which we dub ‘millinovae’ because their optical luminosities are approximately a thousand times fainter than those of ordinary classical novae.
@@ 1 Introduction
A classical nova eruption occurs on the surface of a white dwarf that is accreting matter from its non-degenerate companion via an accretion disk. A layer of fresh, usually hydrogen-rich gas accumulates on the white dwarf surface and becomes degenerate. The gas is heated, and when hydrogen ultimately ignites, a thermonuclear runaway ejects about of matter from the system with velocities from hundreds to thousands of km s-1(Bode & Evans, 2008). As the ejecta expands and becomes optically thin, the exposed white dwarf surface is observed to emit supersoft X-rays due to stable hydrogen burning of the leftover gas (Osborne, 2015; Page & Shaw, 2022). The maximum X-ray luminosities of classical novae are typically erg s-1 (e.g., Henze et al., 2014). In the optical passbands, they normally reach absolute magnitudes from to (Bode & Evans, 2008).
@@
The accretion disk surrounding the white dwarf is subject to a thermal instability that can lead to dwarf nova outbursts. They occur when gas in the disk reaches a critical temperature, and its viscosity abruptly changes, increasing the mass accretion rate onto the white dwarf (Lasota, 2001). If the instability develops in the outer regions of the accretion disk (outside-in outbursts), the rise to maximum brightness is rapid. However, the light curves of inside-out outbursts (which start in the inner disk regions) are more symmetric, with a relatively slow rise to the peak (Lasota, 2001). Dwarf novae in outburst reach optical absolute magnitudes from to , depending on their orbital period (longer orbital periods mean larger accretion disks, and hence become brighter; Patterson 2011). The X-ray luminosities of dwarf novae in outburst are typically erg s-1, about six orders of magnitude lower than those of classical novae (Schwope et al., 2024; Rodriguez et al., 2024).
@@
Therefore, the discovery of supersoft X-ray emission from the optical transient ASASSN-16oh (Maccarone et al., 2019) was a surprise. The X-ray emission was consistent with that of a 900,000 K blackbody and a luminosity of about erg s-1, similar to that seen from persistent supersoft X-ray sources and classical novae. However, the optical properties of the transient (light curve shape, its amplitude, peak absolute magnitude, optical spectra) did not match those of classical novae at all.
@@
ASASSN-16oh was first detected in December 2016 by the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al. 2014) as a transient in the Small Magellanic Cloud (SMC). Its archival light curve (upper panel of Figure 1) collected by the Optical Gravitational Lensing Experiment (OGLE; Udalski et al. 2015) revealed that the optical outburst commenced at least four months before the ASAS-SN discovery. The object exhibited irregular variability before (and after) the outburst, with a mean brightness of about and . It peaked at , which corresponds to an absolute magnitude of , assuming the SMC distance modulus of (Graczyk et al., 2020) and reddening of (Skowron et al., 2021). The full optical light curve does not match those of classical novae, which normally rise to maximum within hours to a few days (not months). Additionally, the optical spectra revealed narrow ( km s-1) Balmer and He ii emission lines (Maccarone et al., 2019), unlike the broad features (several hundred to a few thousand km s-1) usually observed in classical novae.
@@
The unusual outburst light curve, narrow optical emission lines, and X-ray spectrum led Maccarone et al. (2019) to propose that the X-rays come from a spreading layer—a belt around the white dwarf’s equator near the inner disk edge in which a large fraction of the total accretion energy is emitted. However, the nature of ASASSN-16oh remains controversial, with alternative models being based on a non-ejecting nova outburst (Hillman et al., 2019; Kato et al., 2020).
@@
Given the unusual and unexplained optical and X-ray properties of ASASSN-16oh, we decided to investigate whether this star is one of a kind or representative of a larger, as yet undiscovered, group. We searched for objects with outburst properties similar to those seen in ASASSN-16oh among the 20-year-long light curves of 76 million stars observed toward the Magellanic Clouds by the OGLE survey (Udalski et al., 2015).
@@ 2 OGLE Photometric Data
The OGLE photometric data have been collected using the 1.3-m Warsaw Telescope located at Las Campanas Observatory, Chile, during the OGLE-III (2001–2009; Udalski 2003) and OGLE-IV (2010–2024; Udalski et al. 2015) phases of the project. Different instruments were used in different phases of the survey. The OGLE-III camera consisted of eight CCD detectors, with a plate scale of arcsec pixel-1. The OGLE-IV camera had the same pixel scale but comprised thirty-two detectors, providing a field of view of 1.4 deg2. The typical exposure time was 150 s. The photometric data were extracted using a custom version of the Difference Image Analysis algorithm (Tomaney & Crotts, 1996; Alard & Lupton, 1998) implemented by Woźniak (2000). The OGLE-III and OGLE-IV data have been reduced using the same reference image (Mróz et al., 2024), enabling us to obtain homogeneous, long-term light curves of all analyzed objects. Photometry was calibrated to the standard - and -band systems. See Udalski (2003) and Udalski et al. (2015) for a detailed description of the observing setup and data calibration.
@@ 3 Results
We started with a list of 72,303 objects (54,996 in the Large Magellanic Cloud (LMC) and 17,307 in the SMC) whose light curves contain at least five consecutive data points magnified with respect to the remaining light curve (see Mróz et al. (2024) for more details about the selection of outbursting events). We required that the light curve contains at least one outburst with an amplitude larger than 1 mag and duration between 10 and 600 days. Subsequently, we cross-matched selected objects with the Gaia Data Release 3 archive (Gaia Collaboration et al., 2023) and removed artifacts due to high-proper-motion stars (those with proper motions 10 mas yr-1), which were the largest source of contamination. The light curves of the remaining objects (7412 in the LMC and 2689 in the SMC) were visually inspected, and we selected the initial sample of outbursts similar to those of ASASSN-16oh.
We vetted all selected objects using additional data, including multi-color photometry, Gaia parallaxes and proper motions, and sky images. We removed objects that could be classified as classical novae, supernovae, and active galactic nuclei, leaving us with a list of 29 objects (22 in the LMC and seven in the SMC). Example light curves are presented in Figure 1 (see also Figures 7 and 8 in Appendix A). All objects show long-duration (weeks to months), symmetrical (triangle-shaped) outbursts with amplitudes in ranging from 1.0 to 3.7 mag. Some outbursts are deceivingly symmetric, which led to them being classified as gravitational microlensing events (OGLE-mNOVA-12 = MACHO-LMC-7 and OGLE-mNOVA-13 = MACHO-LMC-23; Alcock et al. 2000), although such events should be achromatic and nonrepeating if they really are due to microlensing. The list of all our selected objects is presented in Table 1.
Given the unique spectral and X-ray characteristics of ASASSN-16oh, we also decided to start the near-real-time monitoring of these selected objects in September 2023, with the goal of triggering follow-up observations should any enter an outburst state.
@@ 4 Outburst of OGLE-mNOVA-11
The outburst of OGLE-mNOVA-11 started shortly afterwards, between 2023 October 15.3 and 26.3 and reached a maximum of and on 2023 December 6.1 (Figure 2). This corresponds to and , assuming the LMC distance modulus of (Pietrzyński et al., 2019) and reddening (Skowron et al., 2021). The star returned to quiescence (in ) in late February 2024, about 120 days after the outburst onset. The outburst duration was shorter in .
We obtained a set of low-resolution spectra of OGLE-mNOVA-11 from 2023 November 22 to December 10 (Figure 3) with the Robert Stobie Spectrograph (Kobulnicky et al. 2003; Burgh et al. 2003) mounted on the Southern African Large Telescope (SALT), as part of the SALT Large Science Program on Transients (Appendix B). The spectra cover the range 3500–6700 Å and reveal narrow ( km s-1) emission lines (Balmer lines, He ii 4686 Å). The Bowen blend (a complex of C iii and N iii lines in the range Å) is also clearly detected from 2023 December 2 to 10. The emission lines are redshifted with a mean radial velocity of km s-1, close to that of the LMC systemic value ( km s-1; van der Marel et al. 2002). Both the light curve shape and optical spectra are very different from those of classical novae.
| Name | RA | Decl. | Comments | |||||
| OGLE-mNOVA-01 | 0.059 | ASASSN-16oh | ||||||
| OGLE-mNOVA-02 | 0.042 | |||||||
| OGLE-mNOVA-03 | 0.028 | |||||||
| OGLE-mNOVA-04 | 0.046 | |||||||
| OGLE-mNOVA-05 | 0.016 | candidate | ||||||
| OGLE-mNOVA-06 | 0.011 | |||||||
| OGLE-mNOVA-07 | 0.068 | |||||||
| OGLE-mNOVA-08 | 0.128 | |||||||
| OGLE-mNOVA-09 | 0.122 | candidate | ||||||
| OGLE-mNOVA-10 | 0.119 | |||||||
| OGLE-mNOVA-11 | 0.100 | |||||||
| OGLE-mNOVA-12 | 0.083 | MACHO-LMC-7 | ||||||
| OGLE-mNOVA-13 | 0.124 | MACHO-LMC-23 | ||||||
| OGLE-mNOVA-14 | 0.092 | |||||||
| OGLE-mNOVA-15 | 0.183 | candidate | ||||||
| OGLE-mNOVA-16 | 0.072 | |||||||
| OGLE-mNOVA-17 | 0.105 | |||||||
| OGLE-mNOVA-18 | 0.094 | |||||||
| OGLE-mNOVA-19 | 0.102 | |||||||
| OGLE-mNOVA-20 | 0.078 | |||||||
| OGLE-mNOVA-21 | 0.078 | |||||||
| OGLE-mNOVA-22 | 0.139 | |||||||
| OGLE-mNOVA-23 | 0.099 | |||||||
| OGLE-mNOVA-24 | 0.082 | |||||||
| OGLE-mNOVA-25 | 0.060 | |||||||
| OGLE-mNOVA-26 | 0.124 | |||||||
| OGLE-mNOVA-27 | 0.263 | |||||||
| OGLE-mNOVA-28 | 0.141 | |||||||
| OGLE-mNOVA-29 | 0.105 |
Note. — The table provides equatorial coordinates (for the epoch J2000), mean -band brightness and color in quiescence, mean -band brightness and color in outburst, and color excess (Skowron et al., 2021) toward the detected objects.
OGLE-mNOVA-11 was also observed by the Neil Gehrels Swift Observatory (Gehrels et al., 2004) five times from 2023 December 6 to 30 (Appendix C). Swift’s X-Ray Telescope (XRT; Burrows et al. 2005) detected a faint X-ray source at the transient’s position, with a 0.3–10 keV count rate declining from count s-1 (December 6) to count s-1 (December 30), see Figure 2. All the data were combined to create a single X-ray spectrum, which can be fitted with a blackbody of K absorbed by a column density cm-2. If we fix to the LMC value ( cm-2), then the temperature is slightly better constrained to K. Assuming this blackbody model, the observed 0.3–10 keV luminosity is approximately erg s-1, or erg s-1 when corrected for absorption, about an order of magnitude smaller than that of ASASSN-16oh, but still substantially higher than dwarf nova X-ray luminosities.

### v2

# Millinovae: A New Class of Transient Supersoft X-ray Sources without a Classical Nova Eruption

@@ Abstract
Some accreting binary systems containing a white dwarf (such as classical novae or persistent supersoft sources) are seen to emit low-energy X-rays with temperatures of 106 K and luminosities exceeding erg s-1. These X-rays are thought to originate from nuclear burning on the white dwarf surface, either caused by a thermonuclear runaway (classical novae) or a high mass-accretion rate that sustains steady nuclear burning (persistent sources). The discovery of transient supersoft X-rays from ASASSN-16oh challenged these ideas, as no clear signatures of mass ejection indicative of a classical nova eruption were detected, and the origin of these X-rays remains controversial. It was unclear whether this star was one of a kind or representative of a larger, as yet undiscovered, group. Here, we present the discovery of 29 stars located in the direction of the Magellanic Clouds exhibiting long-duration, symmetrical optical outbursts similar to that seen in ASASSN-16oh. We observed one of these objects during an optical outburst and found it to be emitting transient supersoft X-rays, while no signatures of mass ejection (indicative of a classical nova eruption) were detected. We therefore propose that these objects form a homogeneous group of transient supersoft X-ray sources, which we dub “millinovae” because their optical luminosities are approximately a thousand times fainter than those of ordinary classical novae.
@@ 1 Introduction
A classical nova eruption occurs on the surface of a white dwarf that is accreting matter from its nondegenerate companion via an accretion disk. A layer of fresh, usually hydrogen-rich gas accumulates on the white dwarf surface and becomes degenerate. The gas is heated, and when hydrogen ultimately ignites, a thermonuclear runaway ejects about of matter from the system with velocities from hundreds to thousands of km s-1(Bode & Evans, 2008). As the ejecta expands and becomes optically thin, the exposed white dwarf surface is observed to emit supersoft X-rays due to stable hydrogen burning of the leftover gas (Osborne, 2015; Page & Shaw, 2022). The maximum X-ray luminosities of classical novae are typically erg s-1 (e.g., Henze et al., 2014). In the optical passbands, they normally reach absolute magnitudes from to (Bode & Evans, 2008).
@@
The accretion disk surrounding the white dwarf is subject to a thermal instability that can lead to dwarf nova outbursts. They occur when gas in the disk reaches a critical temperature, and its viscosity abruptly changes, increasing the mass accretion rate onto the white dwarf (Lasota, 2001). If the instability develops in the outer regions of the accretion disk (outside-in outbursts), the rise to maximum brightness is rapid. However, the light curves of inside-out outbursts (which start in the inner disk regions) are more symmetric, with a relatively slow rise to the peak (Lasota, 2001). Dwarf novae in outburst reach optical absolute magnitudes from to , depending on their orbital period (longer orbital periods mean larger accretion disks, and hence, they become brighter; Patterson 2011). The X-ray luminosities of dwarf novae in outbursts are typically erg s-1, about six orders of magnitude lower than those of classical novae (Schwope et al., 2024; Rodriguez et al., 2024).
@@
Therefore, the discovery of supersoft X-ray emission from the optical transient ASASSN-16oh (Maccarone et al., 2019) was a surprise. The X-ray emission was consistent with that of a 900,000 K blackbody and a luminosity of about erg s-1, similar to that seen from persistent supersoft X-ray sources and classical novae. However, the optical properties of the transient (light-curve shape, its amplitude, peak absolute magnitude, and optical spectra) did not match those of classical novae at all.
@@
ASASSN-16oh was first detected in 2016 December by the All-Sky Automated Survey for Supernovae (ASASSN; Shappee et al. 2014) as a transient in the Small Magellanic Cloud (SMC). Its archival light curve (upper panel of Figure 1) collected by the Optical Gravitational Lensing Experiment (OGLE; Udalski et al. 2015) revealed that the optical outburst commenced at least four months before the ASASSN discovery. The object exhibited irregular variability before (and after) the outburst, with a mean brightness of about and . It peaked at , which corresponds to , assuming the SMC distance modulus of (Graczyk et al., 2020) and reddening of (Skowron et al., 2021). The full optical light curve does not match those of classical novae, which normally rise to maximum within hours to a few days (not months). Additionally, the optical spectra revealed narrow ( km s-1) Balmer and He ii emission lines (Maccarone et al., 2019), unlike the broad features (several hundred to a few thousand km s-1) usually observed in classical novae.
@@
The unusual outburst light curve, narrow optical emission lines, and X-ray spectrum led Maccarone et al. (2019) to propose that the X-rays come from a spreading layer—a belt around the white dwarf’s equator near the inner disk edge in which a large fraction of the total accretion energy is emitted. However, this interpretation remains controversial, with alternative models being based on a nonejecting nova outburst (Hillman et al., 2019; Kato et al., 2020).
Hillman et al. (2019) proposed that both the optical flux and X-ray emission in ASASSN-16oh originated from a hot white dwarf that underwent a non-mass-ejecting thermonuclear runaway. They argued that such nonejecting nova events are possible if the mass accretion rate onto the white dwarf is sufficiently high (at least a few times ) and claimed that the light curve of ASASSN-16oh is best modeled by a white dwarf accreting at . Nonetheless, their model predicted an asymmetric outburst (with a rapid rise to the peak followed by a slower decline), in stark contrast to observations. Furthermore Kato et al. (2020) later criticized this model, pointing out that Hillman et al. had overestimated the optical brightness by 6 mag.
On the other hand, Kato et al. (2020) proposed that the outburst of ASASSN-16oh was triggered by a sudden mass accretion caused by disk instability. In this model, the optical flux originates from the irradiated accretion disk, and X-rays are coming from the hot hydrogen-burning white dwarf surface. They found that the X-ray light curve of ASASSN-16oh is best explained by a massive () white dwarf model.
@@
Given the unusual and unexplained optical and X-ray properties of ASASSN-16oh, we decided to investigate whether this star is one of a kind or representative of a larger, as yet undiscovered, group. We thus searched for objects with outburst properties similar to those seen in ASASSN-16oh among the 20-year-long light curves of 76 million stars observed toward the Magellanic Clouds by the OGLE survey (Udalski et al., 2015).
@@ 2 OGLE Photometric Data
The OGLE photometric data have been collected using the 1.3 m Warsaw Telescope located at Las Campanas Observatory, Chile, during the OGLE-III (2001–2009; Udalski 2003) and OGLE-IV (2010–2024; Udalski et al. 2015) phases of the project. Different instruments were used in different phases of the survey. The OGLE-III camera consisted of eight CCD detectors, with a plate scale of arcsec pixel-1. The OGLE-IV camera had the same pixel scale but comprised thirty-two detectors, providing a field of view of 1.4 deg2. The typical exposure time was 150 s. The photometric data were extracted using a custom version of the difference image analysis algorithm (Tomaney & Crotts, 1996; Alard & Lupton, 1998) implemented by Woźniak (2000). The OGLE-III and OGLE-IV data have been reduced using the same reference image (Mróz et al., 2024), enabling us to obtain homogeneous, long-term light curves of all analyzed objects. Photometry was calibrated to the standard - and -band systems. See Udalski (2003) and Udalski et al. (2015) for a detailed description of the observing setup and data calibration. All data presented in this paper are available to the astronomical community from https://ftp.astrouw.edu.pl/ogle/ogle4/millinovae/.
@@ 3 Results
We started with a list of 72,303 objects (54,996 in the Large Magellanic Cloud, LMC, and 17,307 in the SMC) whose light curves contain at least five consecutive data points magnified with respect to the remaining light curve (see Mróz et al. 2024 for more details about the selection of outbursting events). We required that the light curve contain at least one outburst with an amplitude larger than 1 mag and duration between 10 and 600 days. Subsequently, we cross-matched selected objects with the Gaia Data Release 3 archive (Gaia Collaboration et al., 2023) and removed artifacts due to high-proper-motion stars (those with proper motions 10 mas yr-1), which were the largest source of contamination. The light curves of the remaining objects (7412 in the LMC and 2689 in the SMC) were visually inspected, and we selected the initial sample of objects with triangle-shaped, symmetrical outbursts similar to those of ASASSN-16oh.
We vetted all selected objects using additional data, including multicolor photometry, Gaia parallaxes and proper motions, and sky images. We removed objects that could be classified as classical novae, supernovae, and active galactic nuclei, leaving us with a list of 29 objects (22 in the LMC and seven in the SMC). Example light curves are presented in Figure 1 (see also Figures 7 and 8 in Appendix A). All objects show long-duration (weeks to months), symmetrical (triangle-shaped) outbursts with amplitudes in ranging from 1.0 to 3.7 mag. Some outbursts are deceivingly symmetric, which led to them being classified as gravitational microlensing events (OGLE-mNOVA-12 = MACHO-LMC-7 and OGLE-mNOVA-13 = MACHO-LMC-23, Alcock et al. 2000; OGLE-mNOVA-25 = OGLE-LMC-02, Wyrzykowski et al. 2009) although such events should be achromatic and nonrepeating if they really are due to microlensing. The list of all our selected objects is presented in Table 1.
Given the unique spectral and X-ray characteristics of ASASSN-16oh, we also decided to start the near-real-time monitoring of these selected objects in 2023 September, with the goal of triggering follow-up observations should any enter an outburst state.
=2.5in {rotatetable*}
| Name | R.A. | Decl. | Comments | ||||||||
| (mag) | (mag) | (mag) | (mag) | (mag) | (day) | (day mag-1) | (mag) | ||||
| OGLE-mNOVA-01 | 0.059 | ASASSN-16oh | |||||||||
| OGLE-mNOVA-02 | 0.042 | … | |||||||||
| OGLE-mNOVA-03 | 0.028 | … | |||||||||
| OGLE-mNOVA-04 | 0.046 | … | |||||||||
| OGLE-mNOVA-05 | 0.016 | candidate | |||||||||
| OGLE-mNOVA-06 | 0.011 | … | |||||||||
| OGLE-mNOVA-07 | 0.068 | … | |||||||||
| OGLE-mNOVA-08 | 0.128 | … | |||||||||
| OGLE-mNOVA-09 | 0.122 | candidate | |||||||||
| OGLE-mNOVA-10 | 0.119 | … | |||||||||
| OGLE-mNOVA-11 | 0.100 | … | |||||||||
| OGLE-mNOVA-12 | 0.083 | MACHO-LMC-7 | |||||||||
| OGLE-mNOVA-13 | 0.124 | MACHO-LMC-23 | |||||||||
| OGLE-mNOVA-14 | 0.092 | … | |||||||||
| OGLE-mNOVA-15 | 0.183 | candidate | |||||||||
| OGLE-mNOVA-16 | 0.072 | … | |||||||||
| OGLE-mNOVA-17 | 0.105 | … | |||||||||
| OGLE-mNOVA-18 | 0.094 | … | |||||||||
| OGLE-mNOVA-19 | 0.102 | … | |||||||||
| OGLE-mNOVA-20 | 0.078 | … | |||||||||
| OGLE-mNOVA-21 | 0.078 | … | |||||||||
| OGLE-mNOVA-22 | 0.139 | … | |||||||||
| OGLE-mNOVA-23 | 0.099 | … | |||||||||
| OGLE-mNOVA-24 | 0.082 | … | |||||||||
| OGLE-mNOVA-25 | 0.060 | OGLE-LMC-02 | |||||||||
| OGLE-mNOVA-26 | 0.124 | … | |||||||||
| OGLE-mNOVA-27 | 0.263 | … | |||||||||
| OGLE-mNOVA-28 | 0.141 | … | |||||||||
| OGLE-mNOVA-29 | 0.105 | … |
Note. — The table provides equatorial coordinates (for the epoch J2000), mean -band brightness and color in quiescence, mean -band brightness and color in outburst, amplitude of the outburst in -band , duration of the highest-amplitude outburst , mean decline rate , and color excess (Skowron et al., 2021) toward the detected objects.
@@ 4 Outburst of OGLE-mNOVA-11
The outburst of OGLE-mNOVA-11 started shortly afterwards, between 2023 October 15.3 and 26.3 and reached a maximum of and on 2023 December 6.1 (Figure 2), an amplitude of almost 3 mag. This corresponds to and , assuming the LMC distance modulus of (Pietrzyński et al., 2019) and reddening (Skowron et al., 2021). The star returned to quiescence (in ) in late 2024 February, about 120 days after the outburst onset. The outburst duration was shorter in .
We obtained a set of low-resolution spectra of OGLE-mNOVA-11 from 2023 November 22 to December 10 (Figure 3) with the Robert Stobie Spectrograph (Kobulnicky et al. 2003; Burgh et al. 2003) mounted on the Southern African Large Telescope (SALT), as part of the SALT Large Science Programme on Transients (Appendix B). The spectra cover the range 3500–6700 Å and reveal narrow ( km s-1) emission lines (Balmer lines, He ii 4686 Å). The Bowen blend (a complex of C iii and N iii lines in the range Å) is also clearly detected from 2023 December 2 to 10. The emission lines are redshifted with a mean radial velocity of km s-1, close to that of the LMC systemic value ( km s-1; van der Marel et al. 2002). Both the light-curve shape and optical spectra are very different from those of classical novae.
OGLE-mNOVA-11 was also observed by the Neil Gehrels Swift Observatory (Gehrels et al., 2004) five times from 2023 December 6 to 30 (Appendix C). Swift’s X-Ray Telescope (XRT; Burrows et al. 2005) detected a faint X-ray source at the transient’s position, with a 0.3–10 keV count rate declining from count s-1 (December 6) to count s-1 (December 30); see Figure 2. All the data were combined to create a single X-ray spectrum, which can be fitted with a blackbody of K absorbed by a column density cm-2. If we fix to the LMC value ( cm-2), then the temperature is slightly better constrained to K. Assuming this blackbody model, the observed LX (0.3–10 keV) erg s-1, or erg s-1 when corrected for absorption, about an order of magnitude smaller than that of ASASSN-16oh, but still substantially higher than dwarf nova X-ray luminosities.